Genetics question:

In a porcine population, the total genetic variance of weight (expressed in Kg) at 180days is equal to 60. The variance due to dominance effects is equal to 10, that due to epistatic effects at 0 and that due to the environment at 70.
Estimate: (a) heritability in the broad sense
B) heritability in the narrow sense
Answer #1

In the above question,

Given: Vp DominantVariance) = 10

Vi EpistaticVariance) = 0

VE(Varianceduetoenvironmentalsources) = 70

Total Genetic Variance (VG) = 60

We know that; VG = VA+VD+VA

where V_{A}= Additive Genetic Variance

therefore, V_{A}= VG - (Vp + V)

VA = 60 – 10 = 50

Total Phenotypic Variance(V_{P})= Vg + V

therefore, Vp = 60 + 70 = 130

Now, Heritability in broad sense H? = VG/VP

so, H = 60/130 = 0.46

And Heritability in narrow sense uz? = VA/VP

so, h2 = 50/130 = 0.38